OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11

hero (hero):

What do you think it should be multiplied by?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

@Hero

hero (hero):

You need to be able to eliminate one of the variables after multiplying. Is multiplying by 2 going to help you do that?

OpenStudy (anonymous):

I honestly dont know what you wouild multiple by? :/

OpenStudy (radar):

First understand the directions provided by the problem, They want you to add and further they want you to multiply the first equation by a number so when added it sums the x's to zero. Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (radar):

O.K now what is the coefficient of the x in the first equation? It is a 2 What is the coefficient of the x in the 2nd equation? It is a -6. What added to a -6 would equal 0 ? -6 + ? =0 Can you answer that?

OpenStudy (anonymous):

6

OpenStudy (radar):

-6 + 6 = 0 Yes, now what do we have to multiply the first equation so that the 2 on the x becomes a 6?

OpenStudy (radar):

Now remember when you multiply that 2 so it becomes a 6, you must also multiply all of those terms in that first equation.

OpenStudy (anonymous):

3

OpenStudy (radar):

Absolutely, so multiplying through by 3 the first equations becomes: 6x + 9y = 24 Now when added to the second equation your x's will add out.

OpenStudy (anonymous):

so the answer would be 3 to the question?

OpenStudy (radar):

6x + 9y = 24 -6x +4y = 11 Yes that is all they wanted, the answer is 3.

OpenStudy (anonymous):

okay can you help me with another

OpenStudy (anonymous):

Solve the system using the addition method. 3x + 4y = -10 5x - 2y = 18 What is the x-value in the solution?

OpenStudy (anonymous):

no my bad i meant this problem Solve the system using the addition method. 3x + 2y = 10 6x + 4y = 14 What is the solution to the system?

OpenStudy (radar):

Since they are only asking you the x value, all you have to do is eliminate the y's. that can easily be done by multiplying the first equation by -2 being sure you multiply all the way through the first equation. (every term by -2). You should get -6x -4y = 20 6x +4y =14 now adding ------------ 0x +0y = 34 Note that the equation "explodes", It is inconsisent, there are no answers.

OpenStudy (radar):

Note that the left hand side the second is twice the first equation, while on the right sider it is not, therefore it is not consistent, and is not solvable.

OpenStudy (radar):

Be sure you have copied the problem correctly.

OpenStudy (radar):

I notice that the second problem, wanted a solution, well the fact of the matter is there is none.

OpenStudy (radar):

Just for your information, if the 2nd equation had been: 6x + 4y = 40 You would still not have a specific solution as the two are "identities" and there are an infinite number of solutions.

OpenStudy (radar):

Good luck with these, I have to run.